Three power boats in a 500 km handicap race leave at 5 hourly intervals. Boat A leaves first and has a speed for the race of 20 km/h. Boat B leaves 5 hours later and travels at an average speed of 25km/h. boat C leaves last, 5 hours after B, and completes the race at a speed of 40 km/h.
a) Draw a graph of each boat's journey on the same set of axes
b)Use your graphs to find the winner or the race
c)Check your answer algebraically
d) Write a short description of what happened to each boat in the race
PLEASE HELP

Answer by mananth(16949)

y= 20x
500=20x
x=25 hours
red line
...
y = 5+25x
500=5+25x
495=25x
x=495/25
x=19.8 hours
green line
...
y= 10+40x
500=10+40x
490=40x
490/40=x
x= 12.25 hours
Blue line
...
y represents the distance travelled
x the time travelled.
the third boat reaches first, The second boat comes second. The first boat last.
...

m.annath@hotmail.ca